### v1

# CURVALID: Geometrically-guided Adversarial Prompt Detection

Abstract
Adversarial prompts capable of jailbreaking large language models (LLMs) and inducing undesirable behaviours pose a significant obstacle to their safe deployment. Current mitigation strategies rely on activating built-in defence mechanisms or fine-tuning the LLMs, but the fundamental distinctions between adversarial and benign prompts are yet to be understood. In this work, we introduce CurvaLID, a novel defense framework that efficiently detects adversarial prompts by leveraging their geometric properties. It is agnostic to the type of LLM, offering a unified detection framework across diverse adversarial prompts and LLM architectures. CurvaLID builds on the geometric analysis of text prompts to uncover their underlying differences. We theoretically extend the concept of curvature via the Whewell equation into an $n$-dimensional word embedding space, enabling us to quantify local geometric properties, including semantic shifts and curvature in the underlying manifolds. Additionally, we employ Local Intrinsic Dimensionality (LID) to capture geometric features of text prompts within adversarial subspaces. Our findings reveal that adversarial prompts differ fundamentally from benign prompts in terms of their geometric characteristics. Our results demonstrate that CurvaLID delivers superior detection and rejection of adversarial queries, paving the way for safer LLM deployment. The source code can be found at https://github.com/Cancanxxx/CurvaLID.
Curvature, Local intrinsic dimension, Large Language Model, Adversarial Defense
1 Introduction
Large Language Models (LLMs) are widely used in real-world applications such as education, customer service, and legal analysis (Yao et al., 2024). However, adversarial prompts exploit vulnerabilities of LLMs to produce unintended and harmful responses (Wallace et al., 2019; Shen et al., 2023; Deng et al., 2023a). Therefore, ensuring their safety against adversarial prompts is essential to prevent harmful outputs, such as biased language, misinformation, or content inciting physical harassment.
Current defences against adversarial prompts mainly rely on prompt engineering or adversarial training. Input perturbation techniques, like IntentionAnalysis (Zhang et al., 2024a) and SmoothLLM (Robey et al., 2023), operate by modifying the prompt and examining whether the altered version can successfully trigger the LLM’s built-in defence mechanisms. The effectiveness of these methods depends on the degree of perturbation and the robustness of the LLM’s built-in defense mechanisms. Meanwhile, adversarial training approaches, which involve fine-tuning the model to resist adversarial inputs, often struggle to scale to larger LLMs. For example, Latent Adversarial Training (LAT) cannot be easily applied to LLMs exceeding 10 billion parameters (Sheshadri et al., 2024). Given that popular LLMs, such as GPT-3 and PaLM 2, have 175 billion and 340 billion parameters respectively, developing scalable defenses for these large models is essential to ensure their reliability and safety (Anil et al., 2023; Brown, 2020). These models are widely used in applications like automated content generation, customer support, and medical diagnostics, where ensuring robustness against adversarial attacks is critical.
Additionally, these existing solutions are inherently tied to the internal architecture of the targeted LLM, limiting their generality. They do not guarantee consistent performance across different adversarial prompts and varying models. Most importantly, the fundamental differences between adversarial and benign prompts remain poorly understood, underscoring the need for a more generalizable and theoretically grounded defence strategy. Understanding these differences is critical, as adversarial prompts exploit specific vulnerabilities that vary across LLMs, making it difficult for current defenses to address them comprehensively. A deeper understanding could enable the development of effective, robust strategies across diverse models and attack scenarios.
In this work, we introduce CurvaLID, a model-agnostic framework designed to address the limitations of current solutions. CurvaLID provides a defense mechanism that operates independently of the internal architecture of targeted LLMs, ensuring its generality across diverse models. It ensures LLMs safety by detecting and filtering out adversarial prompts before they are being used. CurvaLID distinguishes between benign and adversarial prompts by examining their geometric distinctions. Specifically, it characterizes their Local Intrinsic Dimensionality (LID) at the sentence level and introduces a curvature measure for text at the word level, enabling a comprehensive analysis of the geometric properties of text prompts. LID has previously been shown to characterize adversarial subspaces in the vision domain (Ma et al., 2018), and token-level LID estimation has been explored to measure the truthfulness of LLM-generated text (Yin et al., 2024). However, we find that token-level estimation is insufficient for detecting adversarial prompts, as the local neighbourhood often consists of stop words, providing limited semantic information. To address this, we propose measuring LID at the prompt level—referred to as PromptLID—rather than relying on token-level analysis.
Additionally, we develop TextCurv, a theoretical framework for defining curvature in an $n$-dimensional Euclidean space of word embeddings. By extending the curvature concept from the Whewell equation (Whewell, 1849), we prove that the angle between two tangent vectors is equivalent to the difference in their tangential angles. This provides the foundation for analyzing word-level geometry, enabling us to quantify curvature in text embeddings and capture semantic shifts. Through extensive evaluations, we demonstrate that PromptLID effectively quantifies the high-dimensional local subspaces where adversarial prompts reside, while TextCurv captures curvature at the word level. Together, they reveal semantic shifts and localized structural deviations in the text manifold, providing new insights into the fundamental differences between adversarial and benign prompts.
Our main contributions can be summarized as follows:
  • We propose CurvaLID, a model-agnostic detection framework that leverages geometric distinctions between adversarial and benign prompts. By integrating TextCurv and PromptLID, CurvaLID efficiently and effectively detects adversarial prompts, ensuring robustness across various LLMs.
  • We provide theoretical insights to the design of two novel geometric measures leveraged by CurvaLID. TextCurv extends the Whewell equation to $n$-dimensional Euclidean space, enabling the quantification of semantic shifts through word-level curvature. PromptLID analyzes the local intrinsic dimensionality across entire prompts, capturing adversarial subspaces more effectively.
  • CurvaLID successfully detects about 99% of adversarial prompts, outperforming the state-of-the-art defenses on various adversarial prompts by large margins. It is also highly time-efficient, requiring only 0.25 GPU hours of training, whereas existing adversarial training methods have significantly higher computational costs, usually exceeding 100 GPU hours.
2 Related Work
This section reviews prior research on adversarial attacks and defense mechanisms for LLMs, highlighting their objectives, underlying logic, and key characteristics.
Adversarial attacks on LLM. Adversarial attacks on LLMs involve crafted inputs designed to manipulate models into generating harmful content, like offensive language or dangerous instructions (Zou et al., 2023). These attacks range from a single input (zero-shot) to more complex, continuous dialogue scenarios (multi-shot) (Shen et al., 2023; Dong et al., 2023; Wang et al., 2023). This research focuses on zero-shot text prompt attacks, including techniques like text perturbation that adds gibberish or subtly alters input wording and social-engineered prompts that trick LLMs into harmful behavior (Zou et al., 2023; Schwinn et al., 2023; Chu et al., 2024).
Adversarial defenses for LLM. There are three primary defenses against adversarial attacks on LLMs: input preprocessing, prompt engineering, and adversarial training. Input preprocessing perturbs inputs to disrupt adversarial prompts but may also affect benign ones, with its effectiveness varying depending on the extent of the perturbation and the targeted LLM (Cao et al., 2023; Robey et al., 2023; Yung et al., 2024). Prompt engineering enhances the LLM’s self-defensive capabilities by adding prompts that help expose harmful intent, though its effectiveness varies with different models (Zhang et al., 2024a; Phute et al., 2023; Zhang et al., 2024b). Finally, adversarial training strengthens the LLM’s ability to reject harmful prompts by exposing it to adversarial scenarios during training (Xu et al., 2024b; Jain et al., 2023). However, this approach requires fine-tuning the LLM, making its success dependent on the specific model being protected, with most adversarial training methods confined to white-box models. Additionally, these methods often demand significant computational resources, with training times reaching up to 128 GPU hours (Mazeika et al., 2024) or 12 GPU hours (Sheshadri et al., 2024), and generally exhibit varying effectiveness across different adversarial prompts and LLMs (Sheshadri et al., 2024; Ziegler et al., 2022; Ganguli et al., 2022). Note that our paper belongs to the field of adversarial prompt detection. Unlike existing methods such as perplexity filtering, which rely on LLMs for next-token probability, our method operates independently of the type of the LLM (Hu et al., 2023).
3 Background and Terminology
This section provides a brief overview of the mathematical definitions of LID and curvature, as well as how they are estimated and utilized in practice.
3.1 Local Intrinsic Dimension
Local Intrinsic Dimensionality (LID) measures the intrinsic dimensionality of the local neighborhood around a reference sample (Houle, 2017). Compared to Global Intrinsic Dimension (GID) (Tulchinskii et al., 2024; Pope et al., 2021a), which measures the degree of the $d$-dimension of the global manifold of a data subset, LID focuses on the local neighborhood of given points. Thus, LID is particularly useful in analyzing high-dimensional data with varying dimensionality across the dataset.
Definition 3.1 (Local Intrinsic Dimension (LID)).
(Houle, 2017) LID is mathematically defined as:
$\text{LID}_{F}(r)=\frac{r\cdot F^{\prime}(r)}{F(r)}.$
We are interested in a function $F$ that satisfies the conditions of a cumulative distribution function (CDF) and is continuously differentiable at $r$. The local intrinsic dimension at $x$ is in turn defined as the limit, when the radius $r$ tends to zero:
$\operatorname{LID}^{*}_{F}\triangleq\lim_{r\to 0^{+}}\operatorname{LID}_{F}(r)\,.$
We refer to the LID of a function $F$, or of a point $\mathbf{x}$, whose induced distance distribution has $F$ as its CDF. For simplicity, we use the term ‘LID’ to refer to the quantity $\operatorname{LID}^{*}_{F}$.
$\operatorname{LID}^{*}_{F}$ is the theoretical definition, and in practice, this quantity needs to be estimated (Levina & Bickel, 2004; Tempczyk et al., 2022). In machine learning applications, estimation of LID requires a distance measure and a set of reference points to select nearest neighbors (NN). Following prior work (Gong et al., 2019; Ansuini et al., 2019; Pope et al., 2021b; Zhou et al., 2024a; Huang et al., 2024, 2025), we use the Euclidean distance. The representation of a data point, along with the chosen reference points, significantly influences how LID is interpreted. In adversarial prompt detection, the representation and neighborhood definition directly affect the ability to distinguish between clean and adversarial prompts. Among existing estimators, we use the Method of Moments (MoM) (Amsaleg et al., 2015) due to its simplicity.
3.2 Curvature
The intuition of curvature is how quickly a curve changes direction. In geometry, we can visualize curvature through an osculating circle. Curvature can be measured at a given point by fitting a circle to the curve on which the point resides (Kline, 1998). The formal definition is as follows:
Definition 3.2.
[Curvature measured by osculating circle](Kline, 1998)
The osculating circle for some point $P$ on a curve is defined as the circle tangent to $P$ and passing each other point on the curve. We denote the radius of the osculating circle as $R$. Thus, the curvature $\kappa$ is defined as:
$\kappa=\frac{1}{R}.$
For an arbitrary curve, one can extend the concept of curvature to the rate of tangential angular change with respect to arc length, which is known as the Whewell equation (Whewell, 1849). The tangential angular change refers to the change of angle of inclination of the tangent at the given point.
Definition 3.3.
[Curvature by Whewell equation](Whewell, 1849) Let $s$ be the arc length and tangential angle $\phi$ be the angle between the tangent to point $P$ and the x-axis, for a given point $P$ on a curve. The curvature $\kappa$ is defined as:
$\kappa=\frac{d\phi}{ds}.$
Furthermore, in differential geometry, the curvature can be defined as the change of the unit tangent vector with respect to arc length (Shifrin, 2015; O’Neill, 2006).
Definition 3.4.
[Curvature in differential geometry](Shifrin, 2015) Suppose the curve $\alpha$ is parametrized by arc length $s$ and $\mathbf{T}(s)$ is the unit tangent vector to the curve. We define curvature as
$\kappa(s)=\|\mathbf{T}^{\prime}(s)\|=\left\|\frac{d\mathbf{T}}{ds}\right\|.$
Curvature is also defined and utilized in physics. The Frenet-Serret formulas relate curvature to torsion, tangent, normal, and binormal unit vectors (Frenet, 1852). In the Frenet-Serret formulas, the curvature describes the rotational speed along a curve, which is relevant in kinetics and trajectory applications (Huang et al., 2023). It is also used in autonomous driving, robotics, and quantum computing (Hallgarten et al., 2024; Alsing & Cafaro, 2023; Shabana, 2023).
4 Geometric Analysis and CurvaLID Framework
We aim to develop geometric measures that can effectively characterize benign and adversarial prompts at both the prompt and word levels, and utilize these measurements for adversarial prompt detection. We specifically define our problem statement as follows:
Let $\mathcal{D}=\{(x^{i},y^{i})\}_{i=1}^{n}$ be a labeled dataset comprising $n$ i.i.d. samples $x^{i}$, where each sample is associated with a label $y^{i}$. In this context, each input $x^{i}$ represents a text prompt, with the corresponding label $y^{i}\in\{0,1\}$ indicating whether the prompt is benign ($y^{i}=0$), or adversarial ($y^{i}=1$). Let $\mathcal{M}(x)$ denote the geometric measure applied to a prompt $x$, where $\mathcal{M}(x)$ is composed of two complementary components: the prompt-level measure $\text{PromptLID}(x)$ and the word-level measure $\text{TextCurv}(x)$. These measures are then utilized within an adversarial prompt detection algorithm, formalized as a classification problem where the objective is to minimize the empirical error between the ground-truth labels and the predictions:
$\arg\min_{\theta}\mathbb{E}_{(x,y)\in\mathcal{D}}[\ell(h(\mathcal{M}(x)),y)],$
where $\ell(\cdot)$ denotes the cross-entropy loss function, and $h(\mathcal{M}(x))$ is the classifier applied to the geometric measures $\mathcal{M}(x)=(\text{PromptLID}(x),\text{TextCurv}(x))$. Next, we formally define PromptLID and TextCurv, detailing how they explore the geometric properties at the prompt and word levels, respectively. Finally, we provide an overview of CurvaLID, our adversarial prompt detection model.
4.1 PromptLID: LID estimation at the prompt-level
To characterize the prompt-level geometric properties of benign and adversarial prompts, we propose PromptLID, an LID estimation based on prompt representations obtained from a trained CNN. We first train a model $g$ (CNN) to perform a $k$-class classification task, where the goal is to determine which benign dataset a given prompt belongs to. This involves learning a function $g:\mathcal{B}\rightarrow\mathcal{Q}$ to map the input space $\mathcal{B}$ to the label space $\mathcal{Q}$. The label space is defined as $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{k}\}$, where $k$ is the number of types of benign datasets and is equal to the cardinality of $\mathcal{Q}$. Given a benign prompt dataset =$\mathcal{B}=\{(b,q)^{i}\}_{i=1}^{n}$, where $b$ is the benign prompt and $q$ is its corresponding label, the model learns to classify each prompt into its respective dataset. The objective function used is categorical cross-entropy, as the task involves multi-class classification. The representation $\bm{z}_{1}$, derived from the penultimate dense layer, encodes the prompt as a single vector, which is then used to calculate the PromptLID. The PromptLID expands on the MoM estimation (Amsaleg et al., 2015) of LID on the prompts’ representation in $\bm{z}_{1}$.
Definition 4.1 (PromptLID).
The PromptLID of a given prompt $\bm{x}$ is defined as:
$PromptLID=-k\cdot\frac{\mu_{k}}{\mu_{k}-w^{k}},$
where $k$ is a locality hyperparameter representing the number of nearest neighbors, $\mu_{k}$ is the averaged distance from the prompt representation $\bm{z}_{1}$ of $\bm{x}$ to its $k$-nearest neighbors, and $w^{k}$ is the distance to the $k$-th nearest neighbor.
Given that adversarial prompts often manipulate the high-dimensional space of word embeddings to target rarely encountered subspaces (Szegedy, 2013; Ma et al., 2018), they exhibit distinct geometric characteristics. Adversarial prompts are expected to exhibit higher PromptLID as they push the embeddings into regions of the feature space that are less well-defined and more complex than typical benign inputs. As shown in Section 5.2, PromptLID effectively captures this behavior, highlighting its ability to distinguish between benign and adversarial prompts.
4.2 TextCurv: Curvature at the word-level
To characterize the word-level geometric properties of benign and adversarial prompts, we analyze the curvature of word connections. The aim is to have curvature complement PromptLID by analyzing the word-level geometric properties of prompts, effectively identifying nearly all adversarial prompts. Specifically, curvature captures the relationships between words, revealing subtle semantic shifts based on word order, and uncovering local geometric differences between benign and adversarial prompts.
Previous research indicates that CNN activation creates a curved manifold, evidenced by the significantly higher intrinsic dimension estimated by Principal Component Analysis (PCA) compared to the GID estimated by TwoNN on activation data (Abdi & Williams, 2010; Facco et al., 2017; Ansuini et al., 2019). However, curvature differences between benign and adversarial prompts remain unexplored. Thus, we examine these differences in convolution layers as potential classification features.
Our goal is to establish a definition of text curvature based on existing mathematical definitions, with the curvature capturing semantic shifts according to word sequence and the strength of these shifts. Word sequence plays a crucial role in semantic analysis, helping to accurately capture the local geometric properties of prompts. We focus on the representations of prompts in the convolutional layers of the model $g$ as mentioned in Section 4.1, where the prompt data remains unflattened and in stacked lists of vectors at this stage, which can be viewed as word-level representation. Specifically, we extract the representations $\bm{z}_{2}$ and $\bm{z}_{3}$ from these convolutional layers for further analysis. This stage is critical, as it is where feature spaces are curved, according to prior research (Ansuini et al., 2019).
To capture semantic shifts between consecutive words in a prompt, we draw on Whewell’s equation, where the rate of directional change of a curve is represented by the tangential angular change. In NLP, this angular change is connected to the dot-product formula and cosine similarity, which indicate the semantic similarity or difference between two words (Mikolov, 2013; Levy et al., 2015). We assume that this theory also applies to modern word embeddings like GPT-2 and RoBERTa, and therefore, we define the rate of angular change in text curvature accordingly:
Figure 1: Illustrative diagram of CurvaLID, which classifies benign and adversarial prompts using PromptLID and TextCurv.
Definition 4.2.
[Text Curvature: Rate of angular change]
For any two consecutive word embeddings, denoted by $\vec{u}$ and $\vec{v}$, the rate of angular change, $d\theta$, is defined as:
$d\theta=\arccos\left(\frac{\vec{u}\cdot\vec{v}}{\|\vec{u}\|\|\vec{v}\|}\right).$
However, the rate of angular change alone does not fully capture the semantic shift between words, as it overlooks the magnitude of the shift. In differential geometry, curvature is defined by the rate of change in the tangent vector’s direction relative to the change in arc length. When two curves exhibit the same directional change, the curve achieving this change over a shorter arc length has a higher curvature. Similarly, in text curvature, given the same semantic shift as measured by our rate of angular change, the curvature should increase when the semantic change is more substantial.
We focus on word vector magnitudes to capture the degree of semantic shift. Previous research suggests that magnitude reflects the semantic weight carried by each word and the tokenizers’ understanding of that word within its context (Schakel & Wilson, 2015; Reif et al., 2019). For example, common words tend to have smaller magnitudes due to their frequent use and limited semantic significance (Schakel & Wilson, 2015). Instead of summing vector norms to measure distance changes in curvature—which may seem intuitive and consistent with geometric principles—we sum the inverses of the vector norms. This approach is driven by the hypothesis that larger vector norms signify greater semantic importance, meaning that curvature should be inversely proportional to vector norms, capturing the larger semantic shifts between words.
Lemma 4.3.
[Text Curvature] For any two consecutive word embeddings, denoted by $\vec{u}$ and $\vec{v}$, the text curvature, denoted by TextCurv, is defined as:
$\textit{TextCurv}=\frac{d\theta}{\frac{1}{\|\vec{u}\|}+\frac{1}{\|\vec{v}\|}}.$
The rate of angular change in TextCurv is supported by Theorem 4.4, which links the angle $\theta$ to the difference in tangential angles. Further analysis of how word embedding vector norms relate to arc length is detailed in Appendix A.2.
Theorem 4.4.
For two tangent vectors $\vec{u}$ and $\vec{v}$ in an $n$-dimensional Euclidean space, the angle $\theta$ between them is equivalent to the difference in their tangential angles.
Sketch of Proof.
We employ the Gram-Schmidt process to construct an orthonormal basis for the tangent space, enabling us to express $\vec{u}$ and $\vec{v}$ as linear combinations of the basis vectors. In this basis, the angle $\theta$ between the vectors is determined using their inner product. We then compute the tangential angles of $\vec{u}$ and $\vec{v}$ from their projections onto the orthonormal basis. By isolating these tangential angles and subtracting them, we demonstrate that their difference is equivalent to $\theta$. The detailed proof is given in Appendix A.2. ∎
TextCurv captures subtle word-level geometric shifts when adversarial modifications alter a prompt’s semantic structure. Adversarial prompts often cause larger and more erratic curvature shifts as they introduce perturbations that disrupt the normal flow of meaning. By analyzing these shifts, TextCurv helps us identify adversarial inputs that deviate from the expected smoothness of benign prompts.
4.3 CurvaLID: Adversarial Prompt Classification by Curvature and LID
CurvaLID is an adversarial prompt detection method that filters out adversarial prompts before they are input to LLMs, ensuring their security. Since CurvaLID operates independently of LLMs, it provides a unified defensive performance across all LLMs. This differentiates it from existing SOTA defenses like input perturbation, prompt engineering, and adversarial training, which show varying performance across different adversarial prompts and LLMs. Moreover, CurvaLID’s evaluation is straightforward and standardized, avoiding the need for subjective human assessments or reliance on LLM judgments, which can raise robustness concerns (Chen et al., 2024; Raina et al., 2024).
CurvaLID involves four steps (see Figure 1, pseudo code in Appendix A.1). In the first step, we use our trained model $g$ (defined in Section 4.1) to classify different types of benign prompts, deriving the normal feature manifold. This is essential for amplifying the geometric and dimensional distinctions between benign and adversarial prompts in subsequent steps. In the second step, we extract the representations of both benign and adversarial prompts from $\bm{z}_{1}$, $\bm{z}_{2}$, and $\bm{z}_{3}$. Then, we compute PromptLID and TextCurv.
Then in the final step, we train a Multilayer Perceptron (MLP) to classify benign and adversarial prompts using the two mean TextCurv values, and the PromptLID as inputs. The MLP performs binary classification and filters out adversarial prompts before they reach the LLMs.
5 Experiments
This section highlights CurvaLID’s performance in classifying benign and adversarial prompts. We demonstrate the limitations of word-level LID due to stop words and punctuation. Lastly, we analyze TextCurv, emphasizing the difference between benign and adversarial prompts. Unless specified, the general experimental setup is as follows: we use a total of 2,519 testing prompts, comprising 1,200 benign prompts and 1,319 adversarial prompts. For the benign prompts, we randomly sampled 300 prompts from each of the Orca (Lian et al., 2023), MMLU (Hendrycks et al., 2020), AlpacaEval (Li et al., 2023a), and TruthfulQA (TQA) (Lin et al., 2021) datasets. The model $g$ is trained to classify these 4 benign datasets. Similarly, for the adversarial prompts, approximately 300 were randomly sampled from each of SAP (Deng et al., 2023a), DAN (Shen et al., 2023), MathAttack (Zhou et al., 2024b), and GCG (Zou et al., 2023). Detailed hyperparameter settings and configurations for the CNN and MLP are in Appendix A, B.1.4 and B.2.
We have included ablation studies and comparisons to other defenses in Appendix B.2 and B.4 respectively. It is important to note that SOTA defenses, such as input perturbation and adversarial training, are highly dependent on the specific adversarial prompts used and the particular LLMs targeted. In contrast, CurvaLID operates independently of the LLMs, providing consistent defensive performance across various models. This fundamental difference makes direct comparisons challenging. Nonetheless, we have attempted to summarize the performance of these methods and provide a comparative analysis with our method in Appendix B.4.
5.1 Adversarial Prompt Classification by CurvaLID
This section explores CurvaLID’s performance in classifying adversarial prompts, assessing its ability to identify mixed benign and adversarial inputs in real-world conditions. We also compare CurvaLID’s detection capabilities with existing defenses to demonstrate its effectiveness.
5.1.1 Classification on adversarial prompts
In this subsection, we demonstrate CurvaLID’s effectiveness in accurately classifying benign and adversarial prompts, including both English and non-English adversarial prompts. We compared a total of 8 datasets for our experiment, consisting of four benign datasets, namely Orca, MMLU, AlpacaEval, and TQA, and four adversarial datasets, namely SAP, DAN, MathAttack, and GCG. These datasets are the popular and SOTA English prompt datasets for testing the robustness of LLMs. From each dataset, we randomly selected at least 300 prompts, resulting in over 2,500 prompts for our classification experiment. We evaluated the ability of CurvaLID to distinguish between benign and adversarial prompts. The experimental setup used here is consistent with that of the following experiments unless specified. Further details about the datasets are in Appendix B.1.5.
CurvaLID on English adversarial prompts. The experimental results in Table 1 demonstrate CurvaLID’s high performance. The model achieved an overall accuracy of 0.992, with perfect accuracy of 1.00 (i.e., 100%) in identifying adversarial prompts and 0.984 accuracy in identifying benign prompts. Since adversarial prompts are detected before reaching the LLM, a detection accuracy of 1.00 directly translates to an attack success rate of 0, effectively nullifying adversarial attempts.
In addition to evaluating CurvaLID on the four main adversarial datasets, which contain a substantial number of prompts, we extended our experiments to other datasets like RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack (Chao et al., 2023; Andriushchenko et al., 2024; Liao & Sun, 2024; Zeng et al., 2024; Liu et al., 2023; Li et al., 2024). These datasets often include highly similar prompts or variants of those in our main experiments. CurvaLID can effectively identify all adversarial attacks with near 0.99 accuracy. Detailed results are in Appendix B.2.1.
CurvaLID on non-English adversarial prompts. We analyzed CurvaLID’s effectiveness in detecting non-English adversarial prompts by testing it on nine languages from the MultiJail dataset (Deng et al., 2023b). We randomly sampled 300 prompts from each of the 9 languages in the MultiJail dataset and tested them individually against 400 benign prompts, adjusted to avoid class imbalance. We gathered 100 prompts from each of four benign datasets randomly. Notably, retraining with benign prompts in different languages was unnecessary to maintain high accuracy. As shown in Table 2, CurvaLID achieved an overall accuracy and F1 score of 0.994, highlighting its robust ability to detect adversarial prompts across languages.
5.1.2 Analysis on CurvaLID
We conducted an extensive analysis of CurvaLID’s performance, including experiments with reduced training prompt sizes and training without adversarial prompts.
Impact of number of prompts on CurvaLID. We studied the impact of reducing the number of prompts on CurvaLID’s performance. We decreased the number of prompts in each dataset to 150 (from 300). Despite this reduction, CurvaLID maintained an overall accuracy of 0.988 and an F1 score of 0.988. Further details of the experimental results can be found in Appendix B.2.2.
Using CurvaLID in one-class classification problems. We experimented at modifying step 4 of CurvaLID by replacing the supervised MLP with unsupervised outlier detection methods such as the local outlier factor (LOF) (Breunig et al., 2000) or isolation forest (Liu et al., 2008), which do not require training on adversarial prompts. Despite the absence of adversarial examples during training, the LOF and isolation forest methods achieved a comparable detection accuracy of around 0.9. Therefore, our framework is task-independent and can be applied to various problem settings. Further details are provided in Appendix B.2.3.
Table 1: Performance metrics for CurvaLID on benign and adversarial datasets.
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.984 0.992 0.992 0.968 1.000 0.983 0.986 Adv. SAP DAN MathAtk GCG 1.000 1.000 1.000 1.000 1.000
5.2 Analysis of local and global ID
Table 2: Performance metrics for CurvaLID on non-English adversarial datasets.
Adv. Dataset zh it vi ar ko th bn sw jv Avg Benign Acc. 0.975 1.000 1.000 1.000 1.000 1.000 1.000 0.975 1.000 0.994 Adv. Acc. 1.000 0.984 0.984 1.000 1.000 1.000 0.984 1.000 0.984 0.993 Overall Acc. 0.988 0.994 0.994 1.000 1.000 1.000 0.994 0.988 0.994 0.994 F1 Score 0.987 0.994 0.994 1.000 1.000 1.000 0.994 0.987 0.994 0.994
Table 3: Average TextCurv of benign and adversarial datasets across word embedding and CNN layers. Percentages in parentheses represent the increase in TextCurv for adversarial prompts compared to benign prompts.
Word Embedding Conv Layer 1 Conv Layer 2 Benign Adv. Benign Adv. RoBERTa 0.626 0.881 (+40.7%) 0.325 0.446 (+37.2%) GPT-2 0.805 1.11 (+37.9%) 0.389 0.546 (+40.4%) BERT 0.428 0.590 (+37.9%) 0.199 0.264 (+32.7%) XLNet 0.296 0.431 (+45.6%) 0.199 0.264 (+32.7%) DistilBERT 0.386 0.557 (+44.3%) 0.225 0.322 (+43.1%)
This subsection addresses the limitations associated with using token-level LID, for classifying benign and adversarial prompts. Adversarial prompts manipulate rarely encountered areas of the feature space using complex words and irregular combinations (Wallace et al., 2019; Ilyas et al., 2019; Ren et al., 2019). Existing work has shown that adversarial perturbation leads the representation to a subspace with distinct local dimensional properties (high LID) (Ma et al., 2018). Hence, we hypothesize that each word in adversarial prompts could lead to a representation with high LID. By utilizing token-level representations and aggregating the LIDs for each prompt by taking the average, it is possible to detect adversarial prompts.
Effectiveness of token-level LID values in separating benign and adversarial prompts. Our analysis revealed significant overlap in token-level LID values between benign and adversarial prompts. We calculated average word-level LID using RoBERTa embeddings, treating each word as a data point with its neighborhood defined by the prompt. As shown in Figures LABEL:fig:LID_SD and Appendix B.3.3, the average token-level LID across most datasets is around 10, making it difficult to differentiate between benign and adversarial prompts. Additionally, the standard deviations are notably high, with AlpacaEval exceeding 100 and DAN surpassing 200. This variability results in substantial overlap in LID values, reducing the effectiveness of token-level LID for distinguishing between prompts.
Investigating similarity benign and adversarial token-level LIDs. Next, we investigated the underlying reasons for the similarity in average token-level LID between benign and adversarial prompts. We focused on the average of the first 5 nearest-neighbor distances (NN-distance), which, as shown in Figure LABEL:fig:KNNdist, appear similar for both benign and adversarial datasets. We present the top 10 most common nearest-neighbors for different prompts in different datasets in Appendix B.3.2. It shows that common nearest-neighbors in the representation space were mostly stop words and punctuation. This suggests that word-level LID overlooks the sequential order of text and often relies on conjunctions, articles, and punctuation. These findings suggest that word-level LID is ineffective for detecting adversarial prompts.
Effectiveness of token-level LID after removing stop words and punctuation. To assess the impact of stop words and punctuation on LID, we recalculated word-level LID after removing them. As shown in Appendix B.3.3, this adjustment reduces the standard deviation to a maximum of 2.26, yielding more consistent LID values. However, the difference between benign and adversarial prompts’ LID remains insufficient for discrimination.
Analysis on PromptLID. We analyze PromptLID across benign and adversarial datasets. As shown in Figure LABEL:fig:LID_comp, adversarial prompts exhibit a much higher average PromptLID (19.472) compared to benign prompts (3.509), highlighting its effectiveness in distinguishing adversarial prompts. The distribution of PromptLID between benign and adversarial prompts can be found in Appendix B.2.14.
Effectiveness of global intrinsic dimension in adversarial prompt detection. Global intrinsic dimension (GID) is another plausible approach for the word-level representation. It can avoid aggregating the LID for each word by assessing the GID for each word within the prompt and output a single value. We use the MLE-based estimate from Tulchinskii et al. (Tulchinskii et al., 2024). However, as shown in Figure LABEL:fig:GIDvsLength, GID shows no clear distinction between benign and adversarial datasets. Instead, it strongly correlates with prompt length, with Pearson and Spearman correlation coefficients of 0.92 and 0.98, respectively. Even after removing stop words and punctuation, the results were similar, with a Pearson correlation coefficient of 0.9, highlighting the limitations of GID in detecting adversarial prompts.
5.3 Curvature of benign and adversarial prompts
We analyzed the average TextCurv of adversarial prompts in CurvaLID Step 2. To ensure TextCurv’s independence from the embedding model, we conducted curvature analysis using various embeddings, including GPT-2, BERT, XLNet, and DistilBERT (Radford et al., 2019; Devlin, 2018; Yang, 2019; Sanh, 2019). As shown in Table 3, adversarial prompts consistently exhibited at least 30% higher curvature than benign prompts across all embeddings. The TextCurv distribution for benign and adversarial prompts is also shown in Appendix B.2.14. These findings support our hypothesis that words in adversarial prompts are more irregular and complex than in benign prompts.
We also demonstrated that CNN activation significantly amplifies the TextCurv differences between benign and adversarial prompts. Specifically, the mean TextCurv of adversarial prompts is at least 30% higher than that of benign prompts in the representations of both CNN layers. In contrast, when considering only the word embedding, the TextCurv for benign prompts based solely on word embedding is 4.91, compared to 5.42 for adversarial prompts. The difference in TextCurv is much smaller—just 13%, which is less than half of the difference observed in the CNN layers.
6 Conclusion
In this paper, we introduce CurvaLID, an adversarial prompt detection framework that filters out adversarial prompts before they are input to LLMs, ensuring their security. CurvaLID operates independently of LLMs, providing consistent performance across all models. It achieves over 0.99 accuracy and reduces the ASR of the tested adversarial prompts to near zero. CurvaLID leverages PromptLID and TextCurv, which analyze the geometric properties of prompts at the prompt-level and word-level, respectively. These measures address the limitations of word-level LID caused by stop words and punctuation, forming the foundation for CurvaLID’s robust performance in distinguishing benign and adversarial prompts. Future directions include expanding CurvaLID to spam detection, integrating human feedback, and testing with noisy prompts.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Pseudocode and theoretical proofs for CurvaLID
This section includes the pseudocode and supplementary theoretical proofs for CurvaLID.
A.1 Pseudocode for CurvaLID
Algorithm 1 presents the pseudocode for CurvaLID.
Algorithm 1 CurvaLID
Input: Datasets $\mathcal{D}_{b}$, $\mathcal{D}_{a}$ (benign and adversarial prompts)
Step 1: Data Preparation
- Load datasets $\mathcal{D}_{b}$ and $\mathcal{D}_{a}$.
- Compute word embeddings $E_{b}$ and $E_{a}$ for $\mathcal{D}_{b}$ and $\mathcal{D}_{a}$.
Step 2: Preprocessing
- Pad sequences to uniform length $L_{\text{max}}$.
- Standardize embeddings to zero mean and unit variance.
Step 3: Train CNN for Benign Classification
- Train a CNN $\mathcal{G}$ on $\mathcal{D}_{b}$ to extract prompt-level representations ${\bm{z}}_{1}$.
Step 4: Compute PromptLID and TextCurv
- Calculate PromptLID on ${\bm{z}}_{1}$.
- Extract intermediate layer outputs ${\bm{z}}_{2}$, ${\bm{z}}_{3}$, and calculate TextCurv.
Step 5: Train the Detection Model
- Combine PromptLID and TextCurv as features.
- Train an MLP $\mathcal{H}$ for binary classification of benign vs adversarial prompts.
A.2 Theoretical foundations and mathematical justification of TextCurv
We begin by addressing the reference of TextCurv to Whewell’s equation, specifically focusing on how embedding angles relate to tangential angles. Our objective is to demonstrate that the angle between two word embedding vectors corresponds to the difference in their tangential angles (i.e., the numerator in Whewell’s equation). To this end, we prove that for two tangent vectors, $\vec{u}$ and $\vec{v}$, in $n$-dimensional Euclidean space, the angle $\theta$ between them is equivalent to the difference in their tangential angles. The following proof establishes this equivalence, showing that the angular difference between two vectors directly corresponds to the difference in their tangential angles.
Theorem For two tangent vectors $\vec{u}$ and $\vec{v}$ in $n$-dimensional Euclidean space, the angle $\theta$ between them is equivalent to the difference in their tangential angles.
Proof of Theorem
Step 1: Angle in $n$-Dimensional Space
The angle $\theta$ between two vectors $\vec{u},\vec{v}\in\mathbb{R}^{n}$ is defined as:
$\cos\theta=\frac{\vec{u}\cdot\vec{v}}{\|\vec{u}\|\|\vec{v}\|},$
Step 2: Tangential Angles and Plane Reduction
  • Tangential Angles: Tangential angles describe the orientation of vectors within the specific 2D plane they span. These are defined relative to a chosen reference axis in that plane.
  • Plane Spanned by $\vec{u}$ and $\vec{v}$: Any two vectors in $n$-dimensional space span a 2D subspace (a plane). This means the interaction between $\vec{u}$ and $\vec{v}$ (e.g., the angle $\theta$) is fully determined by their projections into this plane.
  • Orthonormal Basis for the Plane: Using the Gram-Schmidt process, construct an orthonormal basis $\{\vec{e}_{1},\vec{e}_{2}\}$:
    - Normalize $\vec{u}$ to define $\vec{e}_{1}$:
    $\vec{e}_{1}=\frac{\vec{u}}{\|\vec{u}\|}.$
    - Define $\vec{e}_{2}$ as orthogonal to $\vec{e}_{1}$ and lying in the same plane:
    $\vec{e}_{2}=\frac{\vec{v}-(\vec{v}\cdot\vec{e}_{1})\vec{e}_{1}}{\|\vec{v}-(\vec{v}\cdot\vec{e}_{1})\vec{e}_{1}\|}.$
Step 3: Expressing $\vec{u}$ and $\vec{v}$ in the Orthonormal Basis
In the orthonormal basis $\{\vec{e}_{1},\vec{e}_{2}\}$:
$\vec{u}=\|\vec{u}\|\vec{e}_{1},$
and:
$\vec{v}=a\vec{e}_{1}+b\vec{e}_{2},$
where:
$a=\vec{v}\cdot\vec{e}_{1},\quad b=\vec{v}\cdot\vec{e}_{2}.$
Step 4: Computing $\cos\theta$
The angle $\theta$ between $\vec{u}$ and $\vec{v}$ is:
$\cos\theta=\frac{\vec{u}\cdot\vec{v}}{\|\vec{u}\|\|\vec{v}\|}.$
Substituting:
$\vec{u}\cdot\vec{v}=\|\vec{u}\|a,\quad\|\vec{v}\|=\sqrt{a^{2}+b^{2}}.$
Thus:
$\cos\theta=\frac{a}{\sqrt{a^{2}+b^{2}}}.$
Step 5: Computing $\sin\theta$
The magnitude of the cross product $\|\vec{u}\times\vec{v}\|$ in the 2D plane is related to $\sin\theta$ by:
$\|\vec{u}\times\vec{v}\|=\|\vec{u}\|\|\vec{v}\|\sin\theta.$
Substituting:
$\|\vec{u}\times\vec{v}\|=\|\vec{u}\||b|.$
Thus:
$\sin\theta=\frac{|b|}{\sqrt{a^{2}+b^{2}}}.$
Step 6: Computing $\tan\theta$
The tangent of $\theta$ is:
$\tan\theta=\frac{\sin\theta}{\cos\theta}.$
Substituting:
$\tan\theta=\frac{\frac{|b|}{\sqrt{a^{2}+b^{2}}}}{\frac{a}{\sqrt{a^{2}+b^{2}}}}.$
Simplify:
$\tan\theta=\frac{|b|}{a}.$
Thus:
$\theta=|\arctan(b/a)|.$
Step 7: Relating $\theta$ to the Tangential Angles
In the 2D plane:
- The tangential angle of $\vec{u}$ relative to $\vec{e}_{1}$ is:
$\alpha_{\vec{u}}=0\quad(\vec{u}\text{ lies entirely along }\vec{e}_{1}).$
- The tangential angle of $\vec{v}$ is:
$\alpha_{\vec{v}}=\arctan(b/a).$
The difference in tangential angles is:
$|\alpha_{\vec{u}}-\alpha_{\vec{v}}|=|\arctan(b/a)|.$
Thus, the geometric angle $\theta$ between $\vec{u}$ and $\vec{v}$ satisfies:
$\theta=|\alpha_{\vec{u}}-\alpha_{\vec{v}}|.$
This completes the proof. ∎
Now we investigate the relationship between word embedding vector norms and the change of arc length. We start with the goal of approximating the arc length $\Delta s$ between two consecutive word embeddings, $\vec{u}$ and $\vec{v}$, in a high-dimensional space. Arc length is classically defined as the integral of the norm of the tangent vector along the curve. For discrete data points, this is approximated as a sum of the Euclidean distances between points.
1. Discrete Approximation of Arc Length
Given consecutive embeddings $\vec{u}$ and $\vec{v}$, the arc length between these points can be approximated as:
$\Delta s=\|\vec{u}-\vec{v}\|.$
However, directly using $\|\vec{u}-\vec{v}\|$ would treat the embeddings purely as geometric points and ignore their semantic significance as encoded by the vector magnitudes.
2. Semantic Weight and Embedding Norms
In NLP, the norm of a word embedding, $\|\vec{u}\|$, encodes the semantic “weight” or importance of a word within its context (Schakel & Wilson, 2015; Reif et al., 2019). Larger norms indicate that the embedding carries more semantic information, while smaller norms suggest less significance.
For two consecutive embeddings, $\vec{u}$ and $\vec{v}$, their combined semantic importance is proportional to their norms:
$\text{Semantic Importance}\propto\|\vec{u}\|+\|\vec{v}\|.$
3. Inverse Proportionality and Arc Length
To align with the geometric principle in Whewell’s equation that relates curvature ($\kappa$) and arc length ($\Delta s$) as:
$\kappa\propto\frac{1}{\Delta s},$
we posit that arc length ($\Delta s$) should decrease when the semantic importance ($\|\vec{u}\|+\|\vec{v}\|$) increases.
This motivates the choice of the inverse relationship:
$\Delta s\propto\frac{1}{\|\vec{u}\|+\|\vec{v}\|}.$
4. Sum of Inverse Norms as Arc Length
While $\|\vec{u}\|+\|\vec{v}\|$ represents the combined semantic importance of two embeddings, directly using it in the denominator would contradict the inverse proportionality between $\Delta s$ and $\kappa$. Instead, we take the inverse of the norms individually, which ensures the arc length is smaller for larger semantic weights.
Thus, the arc length approximation becomes:
$\Delta s\propto\frac{1}{\|\vec{u}\|}+\frac{1}{\|\vec{v}\|}.$
The reasoning is that embeddings with larger norms (higher semantic significance) should have smaller contributions to the overall arc length, reflecting the sharper semantic transitions between significant words.
Appendix B Supplementary information for experiments
The appendix is organized into four sections. B.1 provides supplementary information about CurvaLID. B.2 focuses on ablation studies, presenting experiments to analyze various aspects of CurvaLID’s performance. B.3 presents supplementary information on LID analysis. Finally, B.4 evaluates the performance of other SOTA defenses, comparing them to CurvaLID. All experiments were conducted on a system with a single Nvidia H100 GPU, 8 CPU cores, and 128 GB of RAM.
B.1 Supplementary information about CurvaLID
This section includes time and space complexity of CurvaLID, and also the experimental settings for CurvaLID analysis.
B.1.1 Time and space complexity of CurvaLID
The time complexity for PromptLID is $\mathcal{O}(np)$, where $n$ is the number of prompts and $p$ is the dimensionality of the prompt embeddings. In our implementation, we use the representation layer of the CNN to obtain the embeddings. For TextCurv, the time complexity is $\mathcal{O}(nmd)$, where $m$ is the number of words in a prompt and $d$ is the word embedding dimensionality. In our case, we use RoBERTa embeddings for word representations. Therefore, the overall time complexity is $\mathcal{O}(n(p+md))$.
The space complexity for PromptLID is $\mathcal{O}(np)$, as we store n prompts, each with $p$ dimensions. For TextCurv, the space complexity is $\mathcal{O}(nz)$, where $z$ represents the dimensionality of the trained CNN layers used in our computations. Consequently, the space complexity is $\mathcal{O}(n(p+z))$.
B.1.2 CNN hyperparameter selection
We conducted a preliminary study for the CNN hyperparameter selection to determine the optimal architecture based on overall CurvaLID detection accuracy, training time, and stability. Stability was evaluated by measuring the average CurvaLID accuracy across 10 random seeds. The study involved experimenting with the following parameters:
Number of convolutional layers: Tested configurations with 1 to 5 layers.
Activation functions: Evaluated ReLU, ELU, tanh, sigmoid, and softplus.
Kernel sizes: Tested kernel sizes ranging from 2 to 5.
Number of parameters: Adjusted the dense layer sizes to 64, 128, and 256 units.
Epochs: Tested training with 10, 20, 30, 40, 50 epochs.
Batch sizes: Evaluated sizes of 16, 32, 64, and 128.
Optimizers: Compared Adam and SGD.
The experimental results are shown in Table 4. We observe that the parameter tuning of the CNN does not significantly impact the overall detection accuracy, as the accuracy generally fluctuates around 0.98 to 0.99.
Table 4: Performance metrics for CNN hyperparameter selection.
Hyperparameter Overall Accuracy Time (min) No. of Conv. Layers 1 0.962 13.2 2 0.992 14.6 3 0.992 14.6 4 0.974 15.8 5 0.993 15.7 Activation Function ReLU 0.992 14.8 ELU 0.993 15.1 tanh 0.977 15.1 Sigmoid 0.975 14.5 Softplus 0.991 14.8 Kernel Size 2 0.942 14.8 3 0.992 14.9 4 0.988 14.8 5 0.982 15.0 Dense Layer Size 64 0.958 14.8 128 0.990 15.2 256 0.992 15.6 Epochs 10 0.943 14.2 20 0.991 14.9 30 0.989 15.3 40 0.990 15.5 50 0.988 15.3 Batch Size 16 0.990 15.6 32 0.991 14.8 64 0.988 14.7 128 0.982 14.2 Optimizer Adam 0.990 15.2 SGD 0.981 15.8
B.1.3 Detailed parameters and specifics of the CNN architecture for classifying benign prompts in CurvaLID step 1
The CNN architecture consists of an input layer and two 1D convolutional layers. The first Conv1D layer applies 32 filters with a kernel size of 3 and a ReLU activation function, while the second Conv1D layer increases the number of filters to 64, again using a kernel size of 3 and ReLU activation. The output from the convolutional layers is flattened before passing through a fully connected layer with 128 units and ReLU activation. Finally, the network includes an output layer with four units and a softmax activation to classify the input into four distinct categories: Orca, MMLU, AlphEval, and TQA. The model is compiled using the Adam optimizer, categorical cross-entropy loss, and accuracy as the evaluation metric. Training is conducted over 20 epochs with a batch size of 32 and a validation split of 20%.
B.1.4 Detailed parameters and specifics of the MLP architecture for classifying benign and adversarial prompts in CurvaLID step 4
The MLP architecture consists of two fully connected layers and an output layer. The first layer contains 256 neurons with ReLU activation, followed by a batch normalization and dropout layer with a rate of 0.5 to prevent overfitting. A second layer, with 128 neurons and ReLU activation, is followed by another batch normalization and dropout layer. The final output layer uses softmax activation with two units corresponding to the binary classification of benign and adversarial prompts. The model is compiled using the Adam optimizer, a learning rate of 0.001, and categorical cross-entropy as the loss function, and it is trained over 150 epochs with early stopping to prevent overfitting.
B.1.5 Experimental settings for Section 5.1
We gathered 2519 prompts in total, with 1200 benign prompts and 1319 adversarial prompts. For benign prompts, we randomly sampled 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA. In addition, we gathered 320 SAP200 prompts, which are 40 randomly selected prompts from 8 adversarial goals. For DAN, we randomly sampled 350 prompts from the adversarial prompts uploaded on their Github, roughly half of their total number of prompts. For MathAttack, we used all 300 adversarial prompts uploaded to their Github. For GCG, we used their default parameter setting: learning rate = 0.01, batch size = 512, top-k = 256, and temperature = 1, and created a universal adversarial suffix. The targetted model is Vicuna-7B-v1.5. We used all 349 adversarial behaviors as listed in their Github.
B.2 Ablation studies
This section presents ablation studies for CurvaLID, including its performance across various adversarial prompts and its effectiveness under different training conditions and parameter settings.
B.2.1 Performance metrics for CurvaLID in RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack
Table 5 shows the performance metrics for CurvaLID in PAIR, RandomSearch, AmpleGCG, Persuassive Attack, AutoDAN, and DrAttack. Note that due to the abundance of each dataset, we are testing these adversarial datasets against benign datasets individually, i.e., four benign datasets against each adversarial dataset.
We obtained 100 prompts from each of the four benign datasets. For PAIR, we sampled 171 adversarial prompts by generating PAIR adversarial prompts targeting Palm2, GPT4, Vicuna-7B, and Llama-2-13b-chat-hf. This is because PAIR’s code focuses on 50 adversarial goals for each LLM, and PAIR does not always find an adversarial prompt for each goal in its default setting. For RandomSearch, we obtain the adversarial prompts from their GitHub directly. We gathered the datasets from attacking vicuna-13B-v1.5, R2D2, Nemotron-4-340b, Llama-2-70b, Llama-2-13B, GPT4_Turbo, GPT3.5_Turbo, and Gemma-7B. We randomly selected 200 unique adversarial prompts, as we observed that some adversarial prompts are the same when attacking different LLMs. For AmpleGCG, we obtained permission to access their adversarial prompts directly. We randomly sampled 200 prompts from the attack to Vicuna-7B on AdvBench. For Persuasive Attack, we included 150 adversarial prompts uploaded by the authors on HuggingFace. For AutoDAN, we included 150 AutoDAN prompts generated by following the setting from their Github, with 50 prompts each targeting Vicuna-7b-v1.5, Llama2-7b, and GPT-4-0613. For DrAttack, we tested 150 adversarial prompts generated by following the setting from their Github, with 50 prompts targeting each of Vicuna-7B, LLaMA-2-7B-Chat, and GPT-3.5-turbo.
Table 5: Performance metrics for CurvaLID in PAIR, RandomSearch, AmpleGCG, Persuasive Attack, AutoDAN, and DrAttack.
Adv. Dataset PAIR RandomSearch AmpleGCG Persuasive Attack AutoDAN DrAttack Avg. Benign Acc. 0.973 1 0.975 0.952 0.973 0.975 0.975 Adv. Acc. 1 1 0.976 1 1 1 0.996 Overall Acc. 0.983 1 0.975 0.962 0.978 0.980 0.986 F1 Score 0.986 1 0.987 0.974 0.986 0.987 0.985
B.2.2 Accuracy of CurvaLID with less data
Table 6 shows the performance of CurvaLID when trained with less data. We training and tested CurvaLID with 150 prompts from each dataset, halving the number of prompts used from the main result. All other parameters remained the same.
We observe that training CurvaLID with less data has a minimal impact on overall detection accuracy, as both the overall accuracy and F1 score only decreased from 0.992 to 0.988. This demonstrates the efficiency of CurvaLID in leveraging geometric features, enabling it to maintain high performance even with limited training data. Such robustness underscores CurvaLID’s potential for deployment in scenarios where access to large, labelled datasets is constrained, making it practical for real-world applications with data scarcity.
Table 6: Classification Accuracy and Performance Metrics for CurvaLID on Benign and Adversarial Datasets with 150 Data from Each Dataset.
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.992 0.988 0.988 0.9565 1.000 1.000 1.000 Adv. SAP DAN MathAtk GCG 0.983 1.000 0.966 1.000 0.969
B.2.3 Performance metrics for CurvaLID with replacing MLP by local outlier factor or isolation forest
The parameters of the local outlier factor is as follows: n_neighbors=30, metric=’chebyshev’, leaf_size=10, and p=1.
For isolation forest, the contamination is set as auto.
We are testing 1900 prompts in total, with 100 prompts randomly sampled from each of SAP, DAN, MathAttack, GCG, PAIR, AmpleGCG and RandomSearch, and 300 prompts from each of Orca, MMLU, AlphacaEval and TruthfulQA. The experimental results are shown in Tables 7 and 8
We observe that the accuracy and F1 score of CurvaLID, when using Local Outlier Factor and Isolation Forest, drop from 0.992 to approximately 0.9. Despite this decrease, it is important to note that this version of CurvaLID operates as a one-class classification model, which inherently simplifies the classification task by focusing on distinguishing a single class. The ability of CurvaLID to maintain a decent performance under these constraints highlights its robustness and adaptability, suggesting its potential for handling future and previously unseen adversarial attacks in dynamic real-world settings.
Table 7: Performance metrics for CurvaLID with replacing MLP by local outlier factor
Metric Benign Adversarial Overall Accuracy 0.953 0.840 0.909 F1 Score 0.927 0.878 0.903
Table 8: Performance metrics for CurvaLID with replacing MLP by isolation forest
Metric Benign Adversarial Overall Accuracy 0.910 0.902 0.907 F1 Score 0.953 0.833 0.903
B.2.4 CurvaLID with PromptLID or TextCurv only
We demonstrate that both PromptLID and TextCurv are crucial for achieving optimal performance in CurvaLID. When using only PromptLID as the input feature, the model achieves an accuracy of 0.95. However, combining PromptLID and TextCurv boosts the accuracy to over 0.99, showcasing the complementary nature of these features. This improvement highlights how TextCurv captures additional geometric properties that PromptLID alone cannot, enabling a more comprehensive distinction between benign and adversarial prompts.
Table 9 illustrated the performance of CurvaLID if we only use LID or TextCurv as our features.
Table 9: Ablation study comparing LID and TextCurv for benign and adversarial prompt classification.
PromptLID TextCurv (1st Conv. Layer) TextCurv (2nd Conv. Layer) TextCurv (Both Conv. Layers) Benign Acc. 0.987 0.690 0.738 0.960 Adv. Acc. 0.932 0.833 0.809 0.884 Overall Acc. 0.958 0.783 0.783 0.777 F1 Score 0.958 0.781 0.782 0.776
B.2.5 Accuracy of CurvaLID in different embeddings
We tested CurvaLID using different word embeddings, including popular ones like GPT-2, BERT, XLNet, and DistilBERT. The experimental results show that CurvaLID performs similarly with around 0.99 overall accuracy, regardless of the word embedding used. Therefore, CurvaLID’s classification performance is independent of the specific word embedding used, demonstrating its robustness and adaptability for deployment across different LLMs with varying word embedding representations.
Table 10 shows the accuracy of CurvaLID in different embeddings. The experimental result shows that CurvaLID maintains a high classification accuracy under different word embeddings.
Table 10: Performance of CurvaLID with Different Word Embeddings. The table summarizes the classification accuracy and F1 scores for benign and adversarial prompts using various word embeddings.
Word embedding Benign Prompt Accuracy Adv. Prompt Accuracy Overall Accuracy F1 RoBERTa 0.984 1.000 0.992 0.992 GPT-2 0.973 1.000 0.986 0.986 BERT 0.987 1.000 0.994 0.993 XLNet 0.991 0.989 0.990 0.990 DistilBERT 0.970 0.992 0.982 0.980
B.2.6 CurvaLID in defending Vicuna-7B-v1.5
We measured the reduction in ASR after CurvaLID identified and filtered out the adversarial attacks in Vicuna-7B-v1.5, using the same settings specified in the respective original adversarial prompt papers. As shown in Table 11, CurvaLID successfully reduced the ASR of most attacks to zero, outperforming the studied SOTA defenses. The experimental results demonstrate that CurvaLID is highly applicable to real-world LLMs, effectively safeguarding them by detecting adversarial prompts before they are processed. Moreover, CurvaLID outperforms SOTA defenses, further highlighting its effectiveness and reliability.
Table 11: Attack success rates (ASR) in percentage after CurvaLID in vicuna-7b-v1.5.
SAP DAN MathAttack GCG PAIR RandomSearch AmpleGCG Vanilla 69 41 56 95 98 95 97.5 CurvaLID 0 0 0 0 0 0 2.4
B.2.7 Replacement of CNN in CurvaLID Step 1 with Transformer and RNN models
We experimented with replacing the CNN architecture (Step 1 of CurvaLID) with Transformer and RNN models. The experimental results are shown in Table 12. While both Transformer and RNN achieved comparable detection accuracy (0.98 versus CNN’s 0.992), they required almost two to three times longer training times. Hence, employing CNN in Step 1 of CurvaLID proves to be the optimal choice, maintaining high detection accuracy while ensuring relatively low training time. This demonstrates CurvaLID’s computational efficiency and practicality. Details of the Transformer and RNN configurations are provided below.
Transformer: The transformer has an input layer, a multi-head attention layer with 4 heads and a key dimension of 64, followed by layer normalization, a dense layer with 128 units, dropout (rate of 0.1), and a final layer normalization. The model then flattens the output, adds another dense layer with 128 units, and concludes with a softmax output layer for classification into 4 classes.
RNN: The RNN model begins with an input layer, followed by two stacked LSTM layers with 64 units each (the first LSTM layer returns sequences, while the second does not). After the LSTM layers, there is a dense layer with 128 units and a ReLU activation, followed by a softmax output layer for classification into 4 classes. The model is compiled with the Adam optimizer and categorical cross-entropy loss.
Table 12: Comparison of CurvaLID Architectures
Metric CNN (Original Setting) Transformer RNN Detection accuracy 0.992 0.984 0.989 Overall training time (min) 14.58 39.01 25.10
B.2.8 Performance of CurvaLID on adversarial prompts with reordered word sequences
We utilized GPT-4-o to reorder the words in every sentence of the adversarial prompts while preserving their semantic meaning. The experimental results, presented in Table 13, demonstrate that CurvaLID maintains robust performance, achieving an overall accuracy of 0.984 in detecting these adversarial prompts with altered word order. It is important to note, however, that reordering the words in adversarial prompts may potentially disrupt their effectiveness as attacks, as the content and intent of the original prompts could be compromised.
Table 13: Performance metrics for CurvaLID on adversarial prompts with reordered word sequences.
Class Dataset Accuracy Class Acc. Overall Acc. F1 Benign Orca MMLU AlpEval TQA 0.981 0.984 0.984 0.993 0.983 0.973 0.973 Adv. SAP DAN MathAtk GCG 0.987 0.983 0.963 1.000 1.000
B.2.9 Performance of CurvaLID on reordered persuasive social-engineered adversarial prompts
In this section, we evaluate CurvaLID on PAIR, DAN, and Persuasive attacks, all of which are social-engineered persuasive attacks designed to preserve both semantic meaning and adversarial intent while varying structure (Chao et al., 2023; Shen et al., 2023; Zeng et al., 2024). To introduce linguistic variations, we utilized GPT-4-o to reorder the words in each sentence of the adversarial prompts while maintaining their semantic meaning. The experimental setup remains the same as described in B.2.1, with the addition of 300 DAN prompts. The experimental results, presented in Table 14, show that CurvaLID consistently achieved over 96% detection accuracy across all three attack types and maintained a 0% attack success rate on Vicuna. These findings highlight the robustness of our method, even against sophisticated and linguistically varied prompts specifically crafted to bypass defenses. This robustness underscores CurvaLID’s potential for deployment in real-world scenarios, where adversarial prompts are likely to exploit linguistic diversity to evade detection.
Table 14: Performance of CurvaLID on reordered social-engineered attacks
Adversarial attack PAIR DAN Persuasive Attack Benign accuracy 0.951 0.971 0.966 Adversarial accuracy 0.988 1.000 0.962 Overall accuracy 0.962 0.983 0.964 Attack success rate on Vicuna-7B-v1.5 0 0 0
B.2.10 Differences in PromptLID and TextCurv between benign and adversarial prompts under linguistic reordering
We conducted an experiment to investigate the differences in PromptLID and TextCurv between benign and adversarial prompts after reordering the adversarial prompts. To introduce linguistic variations, we utilized GPT-4-o to reorder the words in each sentence of the adversarial prompts while preserving their semantic meaning. For the benign prompts, we tested 100 samples each from Orca, MMLU, AlpacaEval, and TQA datasets. Similarly, for the adversarial prompts, we tested 100 samples each from PAIR, DAN, and Persuasive attacks (Chao et al., 2023; Shen et al., 2023; Zeng et al., 2024). Table 15 highlights the geometric differences between benign and adversarial prompts, demonstrating the effectiveness of our method in capturing these variations. Notably, even after linguistic reordering of the prompts, both PromptLID and TextCurv continue to exhibit significant distinctions between benign and adversarial prompts. This ensures that the performance of CurvaLID remains unaffected by such reordering, further underscoring the robustness and reliability of these geometric measures in differentiating adversarial inputs.
Table 15: Geometric differences in TextCurv and PromptLID between benign and adversarial prompts. The percentages in parentheses indicate the relative increase in adversarial prompts compared to benign prompts, calculated as $(\text{Adversarial}-\text{Benign})/\text{Benign}\times 100$.
Geometric Measures TextCurv@Conv Layer 1 TextCurv@Conv Layer 2 PromptLID@Dense Layer Benign Adversarial Benign Adversarial Benign Adversarial Average Value 0.644 0.813 (+26.3%) 0.341 0.425 (+24.6%) 3.546 18.223 (+413.8%)
B.2.11 CurvaLID with separated benign training and testing data
We conducted an ablation study by training CurvaLID on two benign datasets and testing it on the remaining two. Specifically, we trained CurvaLID using only Orca and MMLU as benign data and evaluated it on AlpacaEval and TQA. The results, shown in Table 16, demonstrate an overall detection accuracy of 0.982, just one percentage point lower than when trained on all four benign datasets. These findings indicate that CurvaLID’s performance remains robust and is not overly optimistic, even when tested on unseen benign datasets. This suggests that the geometric features captured by PromptLID and TextCurv generalize well across different benign datasets, reinforcing the adaptability of CurvaLID in scenarios where access to a comprehensive set of benign data may be limited.
Table 16: CurvaLID with different training and testing benign datasets
Data class Accuracy by dataset Accuracy by class Overall accuracy F1 score Benign AlpacaEval TQA 0.9412 0.982 0.98 0.963 0.9194 Adversarial SAP DAN MathAttack GCG 1 1 1 1 1
B.2.12 CurvaLID with long text length benign prompts
We conducted an additional experiment to evaluate CurvaLID’s performance on benign prompts with longer text lengths. Specifically, we calculated the median text length of benign prompts (106 characters in our experiment), removed all benign prompts with fewer than the median, and reevaluated CurvaLID’s performance as outlined in Section 5.1 (Page 7, Line 351). The results, presented in Table 17 below, show that this adjustment had minimal effect on detection accuracy. The overall accuracy was 0.990, compared to 0.992 when all benign prompts (without filtering by text length) were included. This confirms that text length has minimal impact on CurvaLID’s performance, highlighting its robustness and ability to generalize across prompts of varying lengths. Such adaptability makes CurvaLID particularly well-suited for real-world applications, where input lengths can vary significantly.
Table 17: CurvaLID on benign prompts over 106 characters
Data class Accuracy by dataset Accuracy by class Overall accuracy F1 score Benign Orca MMLU AlpacaEval TQA 0.981 0.990 0.990 0.922 1.000 1.000 1.000 Adversarial SAP DAN MathAttack GCG 1.000 1.000 1.000 1.000 1.000
B.2.13 CurvaLID with non-standard benign prompts
We conducted an experiment to evaluate CurvaLID’s performance on non-standard benign samples. Specifically, we utilized GPT-4-o to introduce spelling errors by replacing one word in each sentence of all benign prompts with a misspelled variant. The experimental results are presented in Table 18 below.
Table 18: CurvaLID on benign prompts with spelling errors
Data class Accuracy by dataset Accuracy by class Overall accuracy F1 score Benign Orca MMLU AlpacaEval TQA 0.985 0.990 0.990 0.957 0.983 1.000 1.000 Adversarial SAP DAN MathAttack GCG 0.995 0.990 0.990 1.000 1.000
Our findings reveal that the detection accuracy by dataset exhibited minimal changes, and the overall accuracy remained almost identical to the original experiment , which involved benign prompts without spelling errors. These results demonstrate that introducing spelling errors has a negligible impact on CurvaLID’s performance, reaffirming its robustness in handling non-standard text inputs.
B.2.14 Analysis of PromptLID and TextCurv distributions
Figures 5, 6, and 7 illustrate the distributions of PromptLID and TextCurv for benign and adversarial prompts. The experiment setting follows B.1.5 and the word embeeding used is RoBERTa. Figure 5 demonstrates that adversarial prompts exhibit a significantly wider range of PromptLID values compared to benign prompts, with higher average values. This suggests that adversarial prompts tend to reside in more complex and sparse regions of the feature space. Figures 6 and 7 display the distributions of TextCurv across the first and second convolution layers, respectively. In both layers, adversarial prompts show consistently higher curvature values, reflecting their tendency to cause greater geometric distortions at the word level. These results highlight the ability of PromptLID and TextCurv to distinguish adversarial prompts based on their unique geometric properties, reinforcing their utility in adversarial prompt detection.
Figure 5: Distribution of PromptLID values for benign (blue) and adversarial (red) prompts, showing the number of data points across different PromptLID ranges.
Figure 6: Distribution of TextCurv values in the first convolution layer for benign (blue) and adversarial (red) prompts, indicating the number of data points for each TextCurv range.
Figure 7: Distribution of TextCurv values in the second convolution layer for benign (blue) and adversarial (red) prompts, indicating the number of data points for each TextCurv range.
B.3 LID analysis
This section includes supplementary information on LID analysis.
B.3.1 LID Estimation using Method of Moments
This section provides the pseudo code for estimating LID by the Method of Moments, see Algorithm B.3.1.
Algorithm 2 LID Estimation using Method of Moments
Input: Dataset, Reference points, Number of neighbors $k$
For each data point in Dataset:
- Compute pairwise distances $r$ between the data point and all points in Reference.
- Sort distances in ascending order and store them as $a$.
- Compute the mean of the first $k-1$ nearest distances:
$m=\frac{1}{k-1}\sum_{i=1}^{k-1}a_{i}.$
- Estimate LID for the data point:
$\text{LID}=\frac{m}{a_{k}-m}.$
Output: LID values for all data points.
B.3.2 Top 10 most common nearest-neighbors for different prompts in different datasets
Table 19 shows the top 10 most common nearest-neighbors for different prompts in different datasets. The results reveal that the common nearest neighbors in the representation space are predominantly stop words and punctuation. This indicates that word-level LID fails to account for the sequential structure of text and relies on conjunctions, articles, and punctuation. Consequently, these findings highlight the limitations of word-level LID in effectively detecting adversarial prompts.
Table 19: Top 10 most common nearest neighbors for each dataset. The angle brackets (<>) are used to specify punctuation and newline characters in the tokenizer. The visible space symbol (␣) represents a space preceding a word or punctuation.
Dataset Top 10 most common nearest neighbors SAP ␣to, ␣and, ␣the, <.>, ␣a, ␣of, <,>, ␣<”>, ␣that, ␣your DAN <,>, <.>, <newline>, ␣the, ␣and, ␣to, ␣you, ␣will, ␣not, ␣is MathAttack ␣the, <,>, ␣of, ␣he, <.>, ␣to, ␣and, ␣she, ␣is, ␣a GCG ␣text, tto, use, ized, <}>, ␣mar, dt, ␣a, <’>, ␣Guide Orca ␣the, <,>, <.>, <newline>, ␣and, ␣a, ␣of, ␣to, ␣is, ␣in MMLU ␣the, <,>, <.>, ␣of, ␣a, ␣to, ␣that, ␣and, ␣in, ␣was AlpacaEval ␣the, ␣a, ␣to, <,>, ␣of, ␣and, ␣I, <.>, ␣for, <?> TQA ␣the, <?>, ␣a, ␣you, ␣is, ␣of, ␣that, ␣to, ␣if, ␣in
B.3.3 Average LID and standard deviation of prompts with and without stop words and punctuation
Table 20 shows the average LID and standard deviation of prompts with and without stop words and punctuation.
Table 20: Comparison of Average LID and Standard Deviation (SD) of Prompts with and without Stop Words and Punctuation
Data Type Dataset With Stop Words and Punctuation Without Stop Words and Punctuation Avg. LID SD Avg. LID SD Benign Orca 9.77 2.52 6.08 1.91 MMLU 9.23 2.86 5.64 1.76 AlpacaEval 13.14 104.58 4.07 2.26 TQA 6.29 3.40 3.93 0.65 Adversarial SAP 12.10 0.55 7.27 0.46 DAN 24.52 244.72 7.18 1.18 MWP 8.98 1.44 4.14 1.28 GCG 8.35 0.40 6.18 0.41
B.3.4 Top 10 most common Nearest Neighbor words after removing stop words and punctuation
Table 21 shows the top 10 most common Nearest Neighbor words after removing stop words and punctuation. The experimental result demonstrates that word-level LID is insufficient for distinguishing between benign and adversarial prompts, even after the removal of stop words and punctuation.
Table 21: Top 10 Most Common Nearest Neighbor Words for each dataset after removing stop words and punctuation. The symbol ␣represents a space preceding a word or punctuation.
Dataset Top 10 most common nearest neighbors SAP ␣Remember, ␣write, ␣goal, ␣mission, ␣act, ␣suicide, ␣Use, ␣phrases, ␣use, ␣refer DAN PT, ␣D, G, AN, ␣Chat, ␣answer, ␣AI, ␣responses, ␣response, ␣respond MathAttack ␣many, ␣much, ␣would, ␣apples, ␣money, ␣20, ␣sold, ␣5, ␣bought, ␣day GCG use, ␣mar, ␣text, dt, end, ate, c, package, ized, ␣t Orca ␣answer, à, º, ␣question, à¦, ␣one, ␣following, ␣said, s, ␣Answer MMLU ␣mortgage, acre, ␣state, ␣contract, ␣deed, ␣would, ␣question, ␣statute, ␣action, s AlpacaEval ␣drinks, ␣gathering, ␣interested, ␣give, ␣time, ␣home, br, ␣trying, ␣dishes, ␣guests TQA ␣say, ks, ␣Oz, ␣established, ␣famous, ␣primed, ␣mirror, es, ␣principle, ␣power
B.4 Performance of other SOTA defenses
We summarize the performance of various defenses across different LLMs, focusing on their ability to reduce the ASR of adversarial prompts and compare this to CurvaLID’s unified performance. CurvaLID outperforms the studied defenses and maintains consistent performance across all LLMs. While we primarily compare four key defenses in this subsection—SmoothLLM (Robey et al., 2023), IntentionAnalysis (Zhang et al., 2024a), RTT3d (Yung et al., 2024), and LAT (Sheshadri et al., 2024)—which are considered SOTA or represent some of the most recent developments, we also experimented with other defenses like Gradient Cuff (Hu et al., 2024), SELFDEFEND (Wang et al., 2024), and SafeDecoding (Xu et al., 2024a).
Given the computational and time constraints associated with replicating results and testing across multiple LLMs, we have cited the performance figures for these defenses from their respective papers. This approach ensures fairness, as different studies report varying results for these defenses when replicating them against different adversarial prompts and across different models. Therefore, we rely on the original reported performances to provide a balanced and consistent comparison. Note that since the results for these defenses are primarily based on English adversarial prompts in their respective papers, our analysis here is focused solely on English prompts.
Defenses based on input perturbation demonstrate mixed results depending on the LLM and the nature of the adversarial attack. For instance, IntentionAnalysis reduces the ASR to between 0.03% and 8.34%, but it struggles with models like Vicuna-7B and MPT-30B-Chat, where the ASR for the SAP attack can reach nearly 20%. Similarly, while SmoothLLM can reduce the ASR to nearly 0% in various LLMs, it fails against PAIR attacks, showing ASRs of 46% in Vicuna-13B and 24% in GPT-4. RTT3d, as the first defense against MathAttack, managed to mitigate 40% of MathAttack in GPT4, but it failed to reduce the ASR to under 10%. LAT achieves near-zero ASR for models like Llama2-7B-Chat and Llama3-8B-Instruct. However, its reliance on a white-box setting and its testing on models with fewer than 10 billion parameters limit its broader applicability.
In the remainder of this subsection, we present the defensive performance of the SOTA defenses. The performance figures for all defenses are directly cited from their original papers due to the computational and time constraints involved in replicating results and testing across various LLMs. This approach ensures consistency and fairness in comparison, as the results reported by different studies often vary when replicating these defenses on different adversarial prompts and models. By relying on the figures from the original sources, we aim to provide an accurate and balanced reflection of each defense’s performance.
We first compared the defenses using a common LLM, LLaMA-2-7B-Chat, as all the studied defenses employ this model. Table 22 presents the ASR comparison between CurvaLID and seven other defenses, showcasing CurvaLID’s exceptional performance in mitigating various adversarial prompts.
Table 22: ASR comparison of CurvaLID with other defenses against different adversarial prompts on LLaMA-2-7B-Chat. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
LLM Defense GCG PAIR DAN AmpleGCG SAP MathAttack RandomSearch LLaMA-2-7B-Chat SmoothLLM 0.1 8 - 0 - - 0 LAT 0.007 0.025 - - - - - Gradient Cuff 0.012 0.23 - - - - - IntentionAnalysis 0 - 0.13 - 0 - - SELFDEFEND 0.09 0.21 0.242 - - - - RTT3d 0.17 0.043 - - - - - SafeDecoding 0 0.04 - - - - - CurvaLID (ours) 0 0 0 0.025 0 0 0
The following are the experimental settings and performances of the seven defenses we studied. Note that it includes various LLMs, namely Vicuna (Chiang et al., 2023), Llama-2 (Touvron et al., 2023), Llama-3 (AI@Meta, 2024), GPT-3.5 (Brown, 2020), GPT-4 (Achiam et al., 2023), PaLM2 (Anil et al., 2023), Claude-1 (Anthropic, 2023a), Claude-2 (Anthropic, 2023b), ChatGLM-6B (Zeng et al., 2022), MPT-30B-Chat (Team, 2023), DeepSeek-67B-Chatand (Bi et al., 2024). It also includes various adversarial prompts, namely, GCG (Zou et al., 2023), PAIR (Chao et al., 2023), DAN (Shen et al., 2023), AmpleGCG (Liao & Sun, 2024), SAP (Deng et al., 2023a), MathAttack (Zhou et al., 2024b), RandomSearch (Andriushchenko et al., 2024), Prefill (Haizelabs, 2023), Many-Shot (Anil et al., 2024), AutoDAN (Liu et al., 2023), TAP (Mehrotra et al., 2023), Jailbroken (Wei et al., 2024), LRL (Yong et al., 2023), DrAttack (Li et al., 2024), Puzzler (Chang et al., 2024), MultiJail (Deng et al., 2023b), DeepInception (Li et al., 2023b), and Template (Yu et al., 2023).
SmoothLLM Detailed experimental settings are referred to (Robey et al., 2023). The experimental results are shown in Table 23.
Table 23: ASR comparison of different LLMs against adversarial prompts under SmoothLLM defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
LLM Adversarial Prompt GCG PAIR RandomSearch AmpleGCG Vicuna-13B-v1.5 0.8 46 44 2 Llama-2-7B-chat 0.1 8 0 0 GPT-3.5 0.8 2 0 0 GPT-4 0.8 24 0 0 PaLM-2 0.9 - - - Claude-1 0.3 - - - Claude-2 0.3 - - -
Latent Adversarial Training Detailed experimental settings are referred to (Sheshadri et al., 2024). The experimental results are shown in Table 24.
Table 24: ASR comparison of different LLMs against adversarial prompts under LAT defense.
LLM Adversarial Prompt PAIR Prefill AutoPrompt GCG Many-Shot Llama2-7B-chat 0.025 0.029 0.006 0.007 0 Llama3-8B-instruct 0.0033 0.0068 0 0.009 0
Gradient Cuff Detailed experimental settings are referred to (Hu et al., 2024). The experimental results are shown in Table 25.
Table 25: ASR comparison of different LLMs against adversarial prompts under Gradient Cuff defense.
LLM Adversarial Prompt GCG AutoDAN PAIR TAP Base64 LRL Llama2-7B-chat 0.012 0.158 0.23 0.05 0.198 0.054 Vicuna-7B-v1.5 0.108 0.508 0.306 0.354 0 0.189
IntentionAnalysis Detailed experimental settings are referred to (Zhang et al., 2024a). The experimental results are shown in Table 26.
Table 26: ASR comparison of different LLMs against adversarial prompts under IntentionAnalysis defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
LLM Adversarial Prompt DAN SAP200 DeepInception GCG AutoDAN ChatGLM-6B 5.48 6.12 0 1 2 LLaMA2-7B-Chat 0.13 0 0 0 0 Vicuna-7B-v1.1 3.42 0.31 0 0 10.5 Vicuna-13B-v1.1 0.94 1.12 0 0 3.5 MPT-30B-Chat 5.38 19.2 4.78 4 - DeepSeek-67B-Chat 3.78 1.56 7.57 2 - GPT-3.5 0.64 0 0 0 -
SELFDEFEND Detailed experimental settings are referred to (Wang et al., 2024). The experimental results are shown in Table 27.
Table 27: ASR comparison of different LLMs against adversarial prompts under SELFDEFEND defense.
LLM Adversarial Prompt DAN GCG AutoDAN PAIR TAP DrAttack Puzzler MultiJail GPT-3.5 0.007 0.18 0.31 0.29 0.02 0.71 0.22 0.203 GPT-4 0.002 0 0.01 0.1 0.08 0.04 0.26 0.012
RTT3d Detailed experimental settings are referred to (Yung et al., 2024). The experimental results are shown in Table 28.
Table 28: ASR comparison of different LLMs against adversarial prompts under RTT3d defense. A dash (-) indicates that experiments were not conducted for this setting in the original paper.
LLM Adversarial Prompt PAIR GCG SAP MathAttack GPT-3.5 - - 0.06 9.8 GPT-4 0.265 - - - Llama-2-13B-Chat 0.043 0.17 - - Vicuna-13B-v1.5 0.26 0.15 - - PaLM-2 0.13 - - -
SafeDecoding Detailed experimental settings are referred to (Xu et al., 2024a). The experimental results are shown in Table 29.
Table 29: ASR comparison of different LLMs against adversarial prompts under SafeDecoding defense.
LLM Adversarial Prompt GCG AutoDAN PAIR DeepInception SAP30 Template Vicuna-7B 0.04 0 0.04 0 0.09 0.05 Llama2-7B-Chat 0 0 0.04 0 0 0